40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Aug 26 17:55:14 2015, Zach, Laser, SiFi, Diode driver problems LD_driver_in_rack.jpgREFL_ThorLabs.jpgREFL_custom.jpg
    Reply  Wed Sep 2 12:53:01 2015, Zach, Laser, SiFi, Diode driver fixed PMC_REFL.jpg
       Reply  Fri Sep 4 01:15:52 2015, Zach, Laser, SiFi, Diode driver fixed (*really* this time) 
          Reply  Sat Sep 5 15:07:55 2015, Zach, Laser, SiFi, Bandwidth and noise with new driver test_PMC_OLTFs_9_4_15.pngPMC_laser_freqnoise_9_4_15.png
Message ID: 1308     Entry time: Wed Aug 26 17:55:14 2015     Reply to this: 1315
Author: Zach 
Type: Laser 
Category: SiFi 
Subject: Diode driver problems 

I now have rack power and a nice LIGO power strip (CRYO:1305)---though I only have a single strip-to-chassis cable for now---and I just received my first laser diode driver from Rich & Co., so I figured I'd test it out. Yesterday, I plugged it in and tested it with both lasers, finding that the output power agreed with the current monitor reading according to the datasheet for each head. In the photo below, the DMM is looking at this current monitor channel, which has a scale of 100 V/A (i.e., the shown current is 100 mA).

Today, I wanted to use the PMC setup I've used before (see CRYO:1205) to remeasure the laser frequency noise and see if it is indeed reduced with the new driver. I had problems.

When retweaking the cavity alignment, I couldn't find a clear resonance by ramping either the PMC PZT or the laser current. I was just using the REFL PD at first, which showed some smallish dips that could have been from a misaligned cavity. Then, I borrowed the 1550nm CCD camera and looked in transmission, where I found a pretty-solid-looking TEM00 mode. Oddly, I couldn't align it much better (judging by the HOMs) and I also couldn't get the cavity to flash normally; it had a very persistent TEM00 spot that was not strongly sensitive to current or PMC PZT positioning. Also, the laser power as measured on the REFL PD seemed to have a strong (10-20%), frequency-modulated oscillation on it.

In an attempt to recover the previous state, I switched back to the ThorLabs LDC201C diode controller, and instantly everything was fine again: there were no oscillations in the laser power, and the cavity flashed as expected when either the laser current or PMC PZT were ramped. In both cases, I used the same ThorLabs TED200C TEC controller and the average DC output laser power was the same with either driver (i.e., ThorLabs vs. custom).

I have no idea what's going on. Below are REFL and cavity length actuation traces with each driver (ThorLabs on the LEFT, custom on the RIGHT). Could the laser be running in a multi-mode state with the custom driver? That would seem to explain why I can't get good cavity fringing with it (see the photos below).

 

 

Side note on pinouts

In trying to diagnose the problem, I noticed something strange. The LM14S2 butterfly mount specifies that, on the DB9 cable coming in from the diode driver, the laser cathode signal should be on pin 7 and the laser anode signal on pin 3. For the ThorLabs laser current drivers (e.g., LDC201C, ITC502, etc.), which should be directly compatible, the pinout has the laser cathode on pin 7 (check) but the anode on pin 8 (wtf?), while pin 3 is supposed to be the "laser diode ground". I think everything still works because these diodes should have their anodes grounded (this needs fact checking, but the LM14S2 mount is supposed to only work with "anode-grounded" diodes, so it seems standard---I couldn't find anything in the RIO documentation to corroborate this).

So, I believe the above shouldn't matter, but that is one big difference between how the current has been applied by the ThorLabs units up until this point and how it is applied with the custom driver now---the ThorLabs units have laser anode on pin 8 and the custom unit has laser anode on pin 3. All that said, Dmass's lasers are using pin 3 via a custom breakout on the table next to the laser heads, and he appears to have no problem with his lasers.

ELOG V3.1.3-